# A Comprehensive Survey of Direct Preference Optimization: Datasets, Theories, Variants, and Applications

### v2

@@ 5.3 More Applications
ConPro (Nguyen et al., 2024) leverages direct preference optimization in helping model to learn severity representation in medical images. Ranking degree of severity in recommendation system requires the learning from preference comparison between various severity classes and normal classes. ConPro achieves new benchmark for classification tasks in severity ranking domain, demonstrating the potential of learning from comparison by DPO algorithm.
Preference learning from human feedback can also be applied to explore text-to-motion generation.  Sheng et al. (2024) fine-tunes MotionGPT (Jiang et al., 2024b) by direct preference optimization with collected preference pairs generated by MotionGPT, which guides the based model to choose the sample with higher policy-assigned probability suggested by the reference policy. Experimental results demonstrate that in this work, DPO algorithm outperforms RLHF in alignment metrics, endowing a more effective approach in text-to-motion tasks.
Softmax-DPO (Chen et al., 2024c) utilizes the DPO’s characteristic of learning from both positive samples and negative samples to instill ranking information to LLMs for the sake of personalized preference recommendation. Learning from the multiple negative samples, Softmax-DPO leverages language models’ reasoning capabilities to predict the next user-preferred item rather than the next-token probability, boosting preference recommendation based on LLMs with direct preference optimization.
DMPO (Bai et al., 2024) aims to bridge the gap between the general data used in pre-training stage and the specific recommendation tasks in application. Different from original DPO fine-tuning’s intention of maximizing the positive sample, DMPO intends to enhance models’ performance on cross-domain recommendations by not simply maximizing the probability of positive samples but also minimizing probability of multiple negative samples. Experimental results indicate that DMPO approach is enhancing models’ performance in recommendation and outperforms traditional sequential approaches as well as LLM-based approaches.
@@ 6 Discussion and Open Challenges
  • First, current automatic methods for dataset construction mostly use proprietary AI models, such as GPT-4, to replace human as preference annotator (Lee et al., 2023). Therefore, the annotation quality is retricted by the capability of AI models, which may lead to inferior alignment preference. Furthermore, as the intelligence of AI models continues to advance, it will become increasingly difficult for even human annotators to accurately distinguish between preferred and dispreferred answers (Burns et al., 2023; Zhao et al., 2024b). Therefore, how to develop an accurate and efficient method to provide large-scale human oversight in the future calls for significant attention.
  • Second, current preference feedback primarily comes from either humans or AI models, relying on a single source and failing to leverage multi-source feedback from real-world environments, tools and specialized models (Liu et al., 2024e; McAleese et al., 2024). Additionally, how to effectively harness such heterogeneous feedback to improve the helpfulness, harmlessness and honesty of target models remains an important area for future research.
  • Third, most preference feedback is coarse-grained (i.e., at the instance level) with some recent efforts exploring fine-grained AI feedback (Yu et al., 2024a; Wu et al., 2024d; Xiao et al., 2024; Lai et al., 2024; Lu et al., 2024b), such as feedback at the sentence or token level. However, this fine-grained feedback has primarily been gathered for specialized tasks like mathematical reasoning or visual question answering, which is insufficient for the development of a general model like OpenAI o1 (OpenAI, 2024b) that can generate long internal chain-of-thoughts. Thus, future research is required to develop accurate and efficient methods for providing such fine-grained feedback.
@@
Owing to its stable training, competitive performance, and computational efficiency, DPO has been widely adopted as an alternative in various applications (Rafailov et al., 2023, 2024a; Ethayarajh et al., 2024; Saeidi et al., 2024). However, DPO also has been shown to exhibit inferior generalization compared to online methods (Xu et al., 2024a; Ivison et al., 2024). To enhance its generalization ability, we propose the following potential directions for future research.
  • First, from the data perspective, existing studies suggest that the inferior performance of DPO stems from the using of offline data (Xu et al., 2024a; Ivison et al., 2024; Wang et al., 2024a). As a result, subsequent studies have introduced various DPO variants to incorporate online data from the policy model into alignment learning  (Yuan et al., 2024d; Xiong et al., 2024; Pang et al., 2024), with the aim of improving the generalization ability of DPO. However, the performance of current methods, such as self-rewarding  (Yuan et al., 2024d) and iterative DPO  (Pang et al., 2024), plateaus after several iterations. Thus, how to further enhance the ability of DPO to exploit additional online data remains a open question.
  • Second, from the perspective of learning objective, DPO align models with human preferences via a preferred answer and a dispreferred answer. However, the degree of preference between different preference pairs is different, which is ignored by the original learning objective of DPO. Consequently, various studies have explored methods to enhance the learning objective’s awareness of the preference difference between the rewards of chosen and rejected responses (Meng et al., 2024; Xu et al., 2024b; Xiao et al., 2024; Azar et al., 2023). However, some issues remain underexplored. For example, enabling alignment awareness of nuanced differences in preference pairs is crucial in scenarios where each token carries significant importance. Additionally, teaching models to reason through DPO alignment poses the challenge of fostering an awareness of intrinsic reasoning abilities in chosen responses, rather than merely following reasoning formats or specific tokens.
  • Lastly, both explicit and implicit reward models predict human preferences using static internal parameters, which inherently impose limitations, such as the inability to access real-time information and the lack of vertical domain knowledge. Consequently, tool-augmented or knowledge-augmented reward modeling has been investigated to construct a more accurate and comprehensive reward signal (Li et al., 2024c; Tian et al., 2023; Lin et al., 2024c). Incorporating these methods into implicit reward modeling, however, remains a promising yet challenging area of research.
@@
  • Mixed-modal models (Lu et al., 2023; Team, 2024; He et al., 2024; Xie et al., 2024) aimed to develop a single unified model capable of handling a wide variety of tasks across vision, language-vision, language, and potentially additional modalities. The core idea behind these models is to homogenize each type of input and output into sequences of discrete vocabulary tokens, thereby unifying multi-modal understanding and generation. DPO interprets autoregressive token generation as a Markov Decision Process (MDP), which makes it inherently suitable for aligning mixed-modal models.
  • The reasoning capabilities of LLMs could be further enhanced through DPO algorithm. The emergence of OpenAI o1 model signifies the new era of shifting the computational resources further into post-training (OpenAI, 2024b). Hiding chain of thought empowers o1 remarable reasoning capabilities, which hints that the pivot of LLM training can be moved to post-training especially reinforcement learning stage after the model reaches a certain scale in pre-training. As a lightweight alignment, DPO algorithm is employed to simplify the complicated multi-model training in the post-training stage while maintaining quality performance. Variant of DPO objective may guide the training model to be rewarded for the more fine-grained chain of thought, reinforcing the learning of correct CoT over the multiple incorrect reasoning steps. The potential of DPO algorithm in enhancing reasoning has not been fully explored in current research.
  • Current works in video generation seek to achieve real-world simulations (Li et al., 2024d). Despite these efforts, challenges persist, such as violations of physical laws and issues related to controllability, including NSFW content and privacy concerns. Recent works by  Prabhudesai et al. (2024) explore backpropagating gradients from trained reward models to a video diffusion model. The direct application of DPO to video generation models remains largely unexplored. Given that DPO is successfully utilized for safety alignment and hallucination mitigation, its integration into video generation models offers promising potential.

### v3

@@ 5.3 More Applications
ConPro Nguyen et al. (2024) leverages direct preference optimization in helping model to learn severity representation in medical images. Ranking degree of severity in recommendation system requires the learning from preference comparison between various severity classes and normal classes. ConPro achieves new benchmark for classification tasks in severity ranking domain, demonstrating the potential of learning from comparison by DPO algorithm.
Preference learning from human feedback can also be applied to explore text-to-motion generation.  Sheng et al. (2024) fine-tunes MotionGPT Jiang et al. (2024b) by direct preference optimization with collected preference pairs generated by MotionGPT, which guides the based model to choose the sample with higher policy-assigned probability suggested by the reference policy. Experimental results demonstrate that in this work, DPO algorithm outperforms RLHF in alignment metrics, endowing a more effective approach in text-to-motion tasks.
Softmax-DPO Chen et al. (2024c) utilizes the DPO’s characteristic of learning from both positive samples and negative samples to instill ranking information to LLMs for the sake of personalized preference recommendation. Learning from the multiple negative samples, Softmax-DPO leverages language models’ reasoning capabilities to predict the next user-preferred item rather than the next-token probability, boosting preference recommendation based on LLMs with direct preference optimization.
DMPO Bai et al. (2024) aims to bridge the gap between the general data used in pre-training stage and the specific recommendation tasks in application. Different from original DPO fine-tuning’s intention of maximizing the positive sample, DMPO intends to enhance models’ performance on cross-domain recommendations by not simply maximizing the probability of positive samples but also minimizing probability of multiple negative samples. Experimental results indicate that DMPO approach is enhancing models’ performance in recommendation and outperforms traditional sequential approaches as well as LLM-based approaches.
@@ 6 Discussion and Open Challenges
  • First, current automatic methods for dataset construction mostly use proprietary AI models, such as GPT-4, to replace human as preference annotator Lee et al. (2023). Therefore, the annotation quality is retricted by the capability of AI models, which may lead to inferior alignment preference. Furthermore, as the intelligence of AI models continues to advance, it will become increasingly difficult for even human annotators to accurately distinguish between preferred and dispreferred answers Burns et al. (2023); Zhao et al. (2024b). Therefore, how to develop an accurate and efficient method to provide large-scale human oversight in the future calls for significant attention.
  • Second, current preference feedback primarily comes from either humans or AI models, relying on a single source and failing to leverage multi-source feedback from real-world environments, tools and specialized models Liu et al. (2024e); McAleese et al. (2024). Additionally, how to effectively harness such heterogeneous feedback to improve the helpfulness, harmlessness and honesty of target models remains an important area for future research.
  • Third, most preference feedback is coarse-grained (i.e., at the instance level) with some recent efforts exploring fine-grained AI feedback Yu et al. (2024a); Wu et al. (2024d); Xiao et al. (2024); Lai et al. (2024); Lu et al. (2024b), such as feedback at the sentence or token level. However, this fine-grained feedback has primarily been gathered for specialized tasks like mathematical reasoning or visual question answering, which is insufficient for the development of a general model like OpenAI o1 OpenAI (2024b) that can generate long internal chain-of-thoughts. Thus, future research is required to develop accurate and efficient methods for providing such fine-grained feedback.
@@
Owing to its stable training, competitive performance, and computational efficiency, DPO has been widely adopted as an alternative in various applications Rafailov et al. (2023, 2024a); Ethayarajh et al. (2024); Saeidi et al. (2024). However, DPO also has been shown to exhibit inferior generalization compared to online methods Xu et al. (2024a); Ivison et al. (2024). To enhance its generalization ability, we propose the following potential directions for future research.
  • First, from the data perspective, existing studies suggest that the inferior performance of DPO stems from the using of offline data Xu et al. (2024a); Ivison et al. (2024); Wang et al. (2024a). As a result, subsequent studies have introduced various DPO variants to incorporate online data from the policy model into alignment learning  Yuan et al. (2024d); Xiong et al. (2024); Pang et al. (2024), with the aim of improving the generalization ability of DPO. However, the performance of current methods, such as self-rewarding  Yuan et al. (2024d) and iterative DPO  Pang et al. (2024), plateaus after several iterations. Thus, how to further enhance the ability of DPO to exploit additional online data remains a open question.
  • Second, from the perspective of learning objective, DPO align models with human preferences via a preferred answer and a dispreferred answer. However, the degree of preference between different preference pairs is different, which is ignored by the original learning objective of DPO. Consequently, various studies have explored methods to enhance the learning objective’s awareness of the preference difference between the rewards of chosen and rejected responses Meng et al. (2024); Xu et al. (2024b); Xiao et al. (2024); Azar et al. (2023). However, some issues remain underexplored. For example, enabling alignment awareness of nuanced differences in preference pairs is crucial in scenarios where each token carries significant importance. Additionally, teaching models to reason through DPO alignment poses the challenge of fostering an awareness of intrinsic reasoning abilities in chosen responses, rather than merely following reasoning formats or specific tokens.
  • Lastly, both explicit and implicit reward models predict human preferences using static internal parameters, which inherently impose limitations, such as the inability to access real-time information and the lack of vertical domain knowledge. Consequently, tool-augmented or knowledge-augmented reward modeling has been investigated to construct a more accurate and comprehensive reward signal Li et al. (2024c); Tian et al. (2023); Lin et al. (2024c). Incorporating these methods into implicit reward modeling, however, remains a promising yet challenging area of research.
@@
  • Mixed-modal models Lu et al. (2023); Team (2024); He et al. (2024); Xie et al. (2024) aimed to develop a single unified model capable of handling a wide variety of tasks across vision, language-vision, language, and potentially additional modalities. The core idea behind these models is to homogenize each type of input and output into sequences of discrete vocabulary tokens, thereby unifying multi-modal understanding and generation. DPO interprets autoregressive token generation as a MDP, which makes it inherently suitable for aligning mixed-modal models.
  • The reasoning capabilities of LLMs could be further enhanced through DPO algorithm. The emergence of OpenAI o1 model signifies the new era of shifting the computational resources further into post-training OpenAI (2024b). Hiding chain of thought empowers o1 remarable reasoning capabilities, which hints that the pivot of LLM training can be moved to post-training especially reinforcement learning stage after the model reaches a certain scale in pre-training. As a lightweight alignment, DPO algorithm is employed to simplify the complicated multi-model training in the post-training stage while maintaining quality performance. Variant of DPO objective may guide the training model to be rewarded for the more fine-grained chain of thought, reinforcing the learning of correct CoT over the multiple incorrect reasoning steps. The potential of DPO algorithm in enhancing reasoning has not been fully explored in current research.
  • Current works in video generation seek to achieve real-world simulations Li et al. (2024d). Despite these efforts, challenges persist, such as violations of physical laws and issues related to controllability, including NSFW content and privacy concerns. Recent works by  Prabhudesai et al. (2024) explore backpropagating gradients from trained reward models to a video diffusion model. The direct application of DPO to video generation models remains largely unexplored. Given that DPO is successfully utilized for safety alignment and hallucination mitigation, its integration into video generation models offers promising potential.